Main INDEX
Monthly INDEX
PREV
NEXT
User name paschke
Log entry time 20:59:02 on April21,2004
Entry number 118786
keyword=SLI not working
SLI not working, it hasn't worked anytime on this shift.
It was reportedly working on day shift. We've requested that MCC
get the help they need to make it work; the OTR can't be used
as it trips the beam too quickly to get an effective measurement, so
without the SLI we're flying pretty much blind.
Run 2608, which was started after our rocky return from the Q1
drop, was taken without a good measurement of energy width. Two
very brief glimpses were seen on the OTR: 270 um at 50 uA, and
190um at 20 uA. It isn't clear what the width will be at
95 uA.